The equation y=-1777x+27,153 can be used to predict the number y of gun deaths in the United States x years after 2000, that is, x=0 corresponds to 2000, x=3 corresponds to 2003, x=6 corresponds to 2006, and so on. Predict the number of gun deaths in 2006 and 2008. In what year will the number of gun deaths be 11,160?
The predicted gun deaths in 2006 will be_____.
The predicted gun deaths in 2008 will be _____.
The predicted number of gun deaths will be 11,160 in the year _____.

Answer by brich(7)

The first question asks what the number of gun deaths will be in the year 2006.
As it shows in the question, x=0 for the year 200[0].
For the year 200[3], x=3
For the year 200[6], x=6
Do you notice that x is always equal to the last number of the year?
I put those numbers in [] so that they would stand out.
So for the number of deaths in 200[6], plug 6 into the equation and solve for y.
y=-1777(6)+27,153
y=-10,662+27,153
y=16,491 deaths

Same thing goes for the second question.
It asks the number of deaths in the year 200[8], so plug 8 into the equation and solve.
y=-1777(8)+27,153
y=-14,216+27153
y=12,927 deaths

Finally it wants to know what year the number of deaths will be 11,160.
So far you have been plugging in the number of years for x, and solving for y, the number of deaths.
This time it gives you y, the number of deaths and it asks you to solve for x, the number of deaths.
So plug 11,160 into the equation for y, and solve for x.
11,610=-1777x+27,153
-15,513=-1777x
x=8.72988 years
I know thats not a nice solid single number but don't freak out.
I wants to know within what year will the gun deaths be 11,610.
Normally you would probably think to round up to 9 but don't do that.
This number of deaths occurs towards the end of the 8th year but still within the 8th year.
So your answer for x will be 8.
don't forget to add that back to the year 2000
And your final answer will be the year 2008

I hope that was helpful
Let me know if you still don't understand